OpenStudy (anonymous):

evaluate the expression : (-0.4)^3=____?

Directrix (directrix):

(-.4) * (-.4) * (-.4) = - .064 (-.4)^3 = -0.064

OpenStudy (anonymous):

(-4/10)^3 =-4^3 / 10^3 -.064